OpenStudy (anonymous):

Let f(x) = x2 + 3x – 4 and g(x) = x + 5. Find f(x) • g(x).

OpenStudy (jiteshmeghwal9):

\[f(x) \times g(x)\]plug the values of f(x) & g(x) as\[(x^2+3x-4) \times (x+5)\]solve this & gt ur answer:)

OpenStudy (anonymous):

@Katiebae Could you please tell us which is your question? Do we have to multiply the functions? Or do we have to composite them?

OpenStudy (jiteshmeghwal9):

\[(x^2+4x-x-4) \times(x+5)\]\[(x(x+4)+(x+4)) \times (x+5)\]\[(x+1)(x+4) \times x+5\]

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

x3 + 3x2 + 16x – 20 correct?